OpenStudy (anonymous):

A parallelogram has the vertices (-1, 2), (4, 4), (2, -1) and (-3, -3). Determine what type of parallelogram. Find the perimeter and area

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

am i supposed to find the distance between (-1,2) (4,4) and (2,-1) (-3,-3)

OpenStudy (anonymous):

A(-1, 2), B(4, 4), C(2, -1) and D(-3, -3). find distances between AB, BC,CD and DA then compare these distances. and find the result.

OpenStudy (anonymous):

what would be it?

OpenStudy (anonymous):

distance between (-1,2) (4,4) = 5.38 and (2,1) (-3,-3) = 6.40

OpenStudy (anonymous):

and others?

OpenStudy (anonymous):

(4,4) (2,-1) = 5.38 and (-3,-3) (-1,2) = 5.38

OpenStudy (anonymous):

@sangya21 what do i do next?

OpenStudy (anonymous):

@sangya21 u there?

OpenStudy (anonymous):

3 sides are equal so getting confused. Let me solve first/

OpenStudy (anonymous):

All sides are equal so it is a square

OpenStudy (anonymous):

but (2,1) and (-3,-3) = 6.40

OpenStudy (anonymous):

oh no i was wrong i didnt put -1 ok so there all equal

OpenStudy (anonymous):

yeah. Now what will be the perimeter and area?

OpenStudy (anonymous):

perimeter of square = 4*5.38 = 21.52 and area = (5.38)^2 = 28.9444

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

Thanks again :)